OpenStudy (anonymous):

True or false: Q(t) =A quantity Q growing exponentially according to the formula Qo5^t has a doubling time of ln2/ln5. Explain

OpenStudy (anonymous):

\[2=5^t\] solve for \(t\) to get the doubling time

OpenStudy (anonymous):

change of base tells you \[2=5^t\iff t=\frac{\ln(2)}{\ln(5)}\] so i guess it is true

OpenStudy (anonymous):

What does Qo stand for in the equation?

OpenStudy (anonymous):

the initial population, i.e. the population when you start counting, at \(t=0\) that is why the subscript it 0 and you write \(Q_0\) instead of \(Q(0)\)